OpenStudy (anonymous):

please help with funtions! Consider the functions f(x) = 2(x − 7) and g(x) = 2x. What is the value of (f + g)(4)? −6 14 2 −2

OpenStudy (anonymous):

(f+g)(4) = f(4) + g(4) = 2(-3) + 2(4) = 8-6 = 2

OpenStudy (anonymous):

thank you